OpenStudy (anonymous):

if f(x)=maximum (cos x,1/2, {sin x}) , o<=x<=2 pi where { . } represents fractional part function then 1. number of points where f(x) is not differentiable

OpenStudy (anonymous):

@ganeshie8 ?

OpenStudy (anonymous):

@ParthKohli ? @mathmale ? @aaronq ? @abb0t ? @babymambo13 ? @CO_oLBoY ? @dan815 ? @eliassaab ? @Freelance1 ? @Hero ? @iambatman ? @jcpd910 ? @Kainui ? @Luigi0210 ?

OpenStudy (anonymous):

I am not studied in trig.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i have made the graph but not sure how to check the differentiability

OpenStudy (dls):

If you have already made the graph then check CORNER points in the graph,that is your answer.

OpenStudy (anonymous):

can u show ur graph so that i can compare mine with that

OpenStudy (anonymous):

@DLS ?

OpenStudy (dls):

\[\Large f(x)=\max (\cos x,1/2, {\sin x}) , o<=x<=2 \pi\] ------------------------------------- This is the graph of cosx in the given domain. http://i.imgur.com/svEYdwT.png This is the graph of {sinx} in the given domain. http://i.imgur.com/DwfBAxf.png And lastly this is 1/2 http://i.imgur.com/ax3XEWz.png ------------------------------------ So now,we will draw the final graph of maximum of the function ------------------------------------- http://i.imgur.com/yUfpJOJ.png (the pink lines)

OpenStudy (dls):

So as you see there are 2 corner points right? (where the function definition changes) so the answer should be 2 (most probably). Do you have the answer?Did your graph match? Did you get it?

OpenStudy (anonymous):

i guess the graph of {sin x } is not correct

OpenStudy (dls):

it will either be 0 or 1 so the lines are correct

OpenStudy (anonymous):

why will it be 0 or 1 it is not a greatest integer function

OpenStudy (anonymous):

{sin pi/6}=0.5

OpenStudy (dls):

oops sorry! give me a second!

OpenStudy (dls):

ill correct it in a second

OpenStudy (dls):

can you draw out your graph ?

OpenStudy (anonymous):

where to draw it ?

OpenStudy (dls):

wait,let me draw mine u check

OpenStudy (dls):

 this should be the final graph,prettier than anything!